When a glass is rubbed with a silk cloth, the rod acquires a charge of magnitude 3.45 nC. (a) Did the glass rod gain or lose electrons/ (b) How many electrons were transferred during the process? (c) What is the change in the mass of the glass rod? (d) What is the change in the mass of the silk cloth?




1
Expert's answer
2022-02-18T08:21:47-0500

Given:

"q=3.45\;\rm nC=3.45*10^{-9}\: C"

"e=1.6*10^{-19}\rm\: C"

"m=\rm 9.1*10^{-31}\: kg"


(a) ย glass rod lose electrons.

(b)

"N=\frac{q}{e}=\frac{3.45*10^{-9}}{1.6*10^{-19}}=2.16*10^{10}"

(c)

"\Delta m=N m\\
=2.16*10^{10}*9.1*10^{-31}=1.96*10^{-20}\:\rm kg"